OpenStudy (anonymous):

The length of a rectangle is 2 ft. less than three times the width. If the area of the rectangle is 40 square feet, find the dimensions.

OpenStudy (anonymous):

Area=length*width length=3w-2 width=w so w(3w-2)=40 once you know the width you can find the length

OpenStudy (anonymous):

Sorry, but can you set that up as a quadratic equation?

OpenStudy (anonymous):

just multiply it out w(3w-2)=40 3w^2-2w=40 3w^2-2w-40=0